Strike Price

Call Price

Put Price

95

7.05

1.81

100

4.11

3.86

105

2.14

6.88

2. Start with $0 in your account.

                                       

a) Short XYZ at $100 and use some of the cash to buy a 95 strike call, investing the rest in zero coupon bonds.

b) Construct profit and payoff graphs for this position (ignoring stock borrow fees).

c) Instead, borrow enough cash to buy the 95 strike put.

d) Construct profit and payoff graphs for this position. (Assume you pay interest r on your loan.)
Answer #1

a) Price of stock at expiration=S

Assume stock borrowing cost =$0

Short XYZ at$100 Long (Buy)

95 Strike Call Price = $7.05

Remaining cash available= (100-7.05)=$ 92.95

Payoff for shorting = 92.95-S

Invest 92.95 in Zero coupon bond

Payoff on long $95 strike call=(S-95)

if S>95 then payoff would look like below

Stock price st expirstio Payoff in cal Interest paid on borrowing 0.000.05875 0.0588 10.00-0.05875 $20.00 -0.05875 30.000.05875 40.00 -0.05875 50.000.05875 60.00 -0.05875 0.000.05875 80.00 -0.05875 30.00 -0.05875 4.34125 110.0014.34125 $120.00 24.34125 $130.00 34.34125 45 Profit and payoff 100.00 140.00 150.00 Stock price t expiro Poyoff in col Payoff or shorting chart Title 0.000.05875 32.35 10.00 -0.0587582.35 20.00 0.05875 72.35 30.00 0.05875 62.35 40.00 0.05875 52.35 50.00 -0.05875 42.35 60.00 0.05875 32.35 70.000.05875 22.35 80.00 0.05875 12.35 30.00-0.05875 2.35 7.05 110.0014.34125 -17.05 $120.00 24.34125 -27.05 $130.00 34.34125 -37.05 5-47.05 55 -57.05 100.00 4.34125 140.00 150.00Stock price at expiration Payoff in call Interest paid on borrowing $0.00 $10.00 $20.00 $30.00 $40.00 $50.00 $60.00 $70.00 $80.00 $90.00 $100.00 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 0.05875 4.94125 110.00 14.94125 120.00 24.94125 4125 44.94125 150.00 54.94125 Profitand payoff 60 50 40 30 20 10 $130.0034.9 $140.00 -Payoff in call --#REF !